OpenStudy (anonymous):

Mardie invested $5,500 in an account that compounded monthly. After 4 years and no further deposits, she had $7,000 in the account. What was the account’s interest rate to the nearest hundredth of a percent?

OpenStudy (anonymous):

7000=5500(1+r/12)^48 ?????????

OpenStudy (dumbcow):

good so far, now try to solve for r divide by 5500 then use logs

OpenStudy (dumbcow):

oops nevermind about the logs, haha take 48th root

OpenStudy (anonymous):

could you show it cuz i get this really huge mumber

OpenStudy (dumbcow):

(7000/5500) = (1+r/12)^48 1.2727 = (1+r/12)^48 (1.2727)^1/48 = 1+ r/12 1.005 = 1 + r/12 .005 = r/12 .06044 = r